OpenStudy (anonymous):

please help write the quadratic function y = x2 ( two is squared ) -3x + 2

OpenStudy (espex):

\[\frac{-b \pm \sqrt{b^2-4ac}}{2a} \] where 'a' is the constant times your 'x^2', b is the constant times your 'x' and c is your lone constant.

OpenStudy (anonymous):

You can find the Vertex (h,k) first and rewrite the equation y=a(x-h)^2+k h=-b/2a k=c-b^2/4a